# Tensor Product Attention Is All You Need

### v3

@@ 3.6 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
TPA QKV Factorization. Let each token’s hidden-state vector be $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, and we follow Section 3.1 to project the entire sequence into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token. The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$. For instance, letting $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, we have:
In practice, we merge all ranks $r$ into a single dimension of the output, reshape, and sum over rank indices; see Section 3.1 for details. The factorization for K and V follows the same pattern.
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
Attention Step and Output Projection. Once we have $\mathbf{Q},\mathbf{K},\mathbf{V}$ factorized per token with RoPE applied on $\mathbf{Q}$ and $\mathbf{K}$, the attention step proceeds for each head $i\in\{1,\dots,h\}$ using (3.4). Finally, concatenating these $h$ heads and then projecting them back using an output weight matrix gives the final attention result, as shown in (3.5).
SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@ 4.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy, 2022), using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). Details on architecture hyperparameters and training hardware are shown in Appendix G.1.
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $6\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 6 and 7 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 8910 and 11 in the appendix.
For the medium-size (353M) models (Tables 2 and 9), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Tables 3), where TPA-KVonly attains the highest average (53.52% zero-shot). And for the XL size (1.5B) models (Table 7), TPA-KVonly attains the highest average (55.03% zero-shot).
Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales. The fully factorized TPA excels on mid-scale models, while TPA-KVonly can rival or surpass it at larger scales. In both cases, factorizing the attention activations shrinks autoregressive KV cache requirements by up to 5×5$\times$5 ×10×10$\times$10 ×, thus enabling much longer context windows under fixed memory budgets. In summary, tensor product attention provides a flexible, memory-efficient alternative to standard multi-head attention, advancing the scalability of modern language models.
@@
5 Related Work
Transformers and Attention. As a sequence-to-sequence architecture Transformer (Vaswani et al., 2017) introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns (Child et al., 2019; Shi et al., 2023; Han et al., 2024; Liang et al., 2024a; Li et al., 2024; Liang et al., 2024b), kernel-based projections (Choromanski et al., 2021), and linearized transformers (Tsai et al., 2019; Katharopoulos et al., 2020; Schlag et al., 2021; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024). To decrease memory usage and circumvent the limitation of memory bandwidth in training, Shazeer (2019) proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle with the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) (Ainslie et al., 2023) divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, Hu et al. (2024) proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
KV Cache Optimization. During the inference time of Transformers, key and value tensors of the previous tokens are repeatedly computed due to their auto-regressive nature. To enhance efficiency, firstly proposed by Ott et al. (2019), these tensors can be cached in memory for future decoding, referred to as the KV cache. However, the KV cache requires additional memory usage and may add to more latencies due to the bandwidth limitation (Adnan et al., 2024). Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024). Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed KV.
@@ 6 Conclusion
Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
We now explore whether it is possible to compute attention scores $\mathbf{Q}\,\mathbf{K}^{\top}$ of Tensor Product Attention (TPA) directly from their factorized forms, thereby reducing floating-point operations.
A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$

### v4

@@ 3.5 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
The specific mechanisms for TPA-based query, key, and value factorization, as well as the attention computation within T6, directly follow the comprehensive descriptions provided in Section 3 (particularly Section 3.1 for factorization, and Equations 3.2 and 3.3 for attention computation). A brief recap of these steps within the T6 model context is provided in Appendix B.
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.5) and Theorem 3.1.
SwiGLU Feed-Forward Network. Following shazeer2020glu ; touvron2023llama , our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@
4 FlashTPA Decoding Algorithm
$\mathbf{B}_{Q}$
$(R_{Q},D)$
$\bm{b}^{K}_{\text{cache}}$
$(M,D)$
$\sum_{D}$
$S^{(1)}$
$(M,R_{Q})$
$\mathbf{A}_{Q}$
$(H,R_{Q})$
$\sum_{R_{Q}}$
$S^{(2)}$
$(M,H)$
$\mathbf{a}^{K}_{\text{cache}}$
$(M,H)$
$\odot$
$\mathcal{L}$
$(H,M)$
Softmax
$\bm{\alpha}$
$(H,M)$
$\mathbf{a}^{V}_{\text{cache}}$
$(M,H)$
$\odot$
$\mathbf{O}^{(A)}$
$(M,H)$
$\bm{b}^{V}_{\text{cache}}$
$(M,E)$
$\sum_{M}$
$\mathbf{O}$
$(H,E)$
Figure 2: Data flow diagram for FlashTPA Decoding. Rectangles represent tensors (blue for inputs, yellow for intermediates, red for final output), circles with $\sum$ or $\odot$ denote Einstein summation contractions or element-wise products respectively, and the green rounded rectangle is the softmax operation. Shapes are shown for a single query ($N=1$) interacting with $M$ cached items. $H$ is the number of heads, $R_{Q}$ is the query rank, and $D,E$ are respective feature dimensions for the $\mathbf{B}_{Q}/\bm{b}^{K}_{\text{cache}}$ and $\bm{b}^{V}_{\text{cache}}$ factors. Scaling factors in softmax are omitted for visual clarity.
For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding. This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values. The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components. This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length.
The detailed definitions of the input factorized components and the step-by-step pseudocode for FlashTPA Decoding are provided in Algorithm 1 (see Appendix D for details). An optimized Triton kernel implementation: Algorithm 3 is also outlined in Appendix D.
This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently. It minimizes memory allocations for large intermediate tensors and reduces the overall computational load compared to materializing full query, key, and value tensors, especially for long sequences. Consequently, TPA is not only memory-efficient due to its smaller KV cache footprint but can also be computationally efficient during inference. The experimental results for FlashTPA decoding time are presented in Section 5.2.
Algorithm 1 FlashTPA Decoding Algorithm
1:$\mathbf{A}_{Q}$: Query A tensor $(B,1,H,R_{Q})$
2:$\mathbf{B}_{Q}$: Query B tensor $(B,1,R_{Q},D)$
3:$\mathbf{a}^{K}_{\text{cache}}$: Cached Key A tensor $(B,M,H)$
4:$\bm{b}^{K}_{\text{cache}}$: Cached Key B tensor $(B,M,D)$
5:$\mathbf{a}^{V}_{\text{cache}}$: Cached Value A tensor $(B,M,H)$
6:$\bm{b}^{V}_{\text{cache}}$: Cached Value B tensor $(B,M,E)$
7:$s_{\text{total}},s_{Q},s_{K},s_{V}$: Scaling factors (with defaults $s_{\text{total}}\leftarrow 1/\sqrt{D}$, $s_{Q}\leftarrow 1/R_{Q}$)
8:$\mathbf{O}$: Output tensor $(B,1,H,E)$
9$\triangleright$ Step 1: Interaction between $\mathbf{B}_{Q}$ and $\bm{b}^{K}_{\text{cache}}$ components
10:$S^{(1)}\leftarrow\text{einsum}(\text{``bnrd, bmd}\rightarrow\text{bnmr''},\mathbf{B}_{Q},\bm{b}^{K}_{\text{cache}})$triangleright Shape: $(B,1,M,R_{Q})$. $S^{(1)}_{b,n,m,r}=\sum_{d}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}$
11$\triangleright$ Step 2: Incorporate $\mathbf{A}_{Q}$ component
12:$S^{(2)}\leftarrow\text{einsum}(\text{``bnhr, bnmr}\rightarrow\text{bnmh''},\mathbf{A}_{Q},S^{(1)})$triangleright Shape: $(B,1,M,H)$. $S^{(2)}_{b,n,m,h}=\sum_{r}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}$
13$\triangleright$ Step 3: Incorporate $\mathbf{a}^{K}_{\text{cache}}$ component to get full logits
14:einsum(“bnmh, $\mathcal{L}\leftarrow\text{einsum}(\text{``bnmh, bmh}\rightarrow\text{bhnm''},S^{(2)},\mathbf{a}^{K}_{\text{cache}})$triangleright Shape: $(B,H,1,M)$. $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}$
15$\triangleright$ Step 4: Apply scaling and Softmax
16:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{M}(\mathcal{L}\cdot s_{Q}\cdot s_{K}\cdot s_{\text{total}})$triangleright Shape: $(B,H,1,M)$. Softmax over cache dimension $M$
17$\triangleright$ Step 5: Compute weighted sum with $\mathbf{a}^{V}_{\text{cache}}$ component
18:$\mathbf{O}^{(A)}\leftarrow\text{einsum}(\text{``bhnm, bmh}\rightarrow\text{bnmh''},\bm{\alpha},\mathbf{a}^{V}_{\text{cache}})$triangleright Shape: $(B,1,M,H)$. $\mathbf{O}^{(A)}_{b,n,m,h}=\bm{\alpha}_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}$
19$\triangleright$ Step 6: Incorporate $\bm{b}^{V}_{\text{cache}}$ component and final scaling
[…]
21:return $\mathbf{O}$
@@ 5.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy2022, ), using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). Details on architecture hyperparameters and training hardware are shown in Appendix J.1.
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix I) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 12 and 13 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 141516 and 17 in the appendix.
For the medium-size (353M) models (Table 2 for 0-shot and Table 15 in appendix for 2-shot), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Table 3), where TPA-KVonly attains the highest average (53.52% zero-shot). For the XL size models (1.5B) (Table 13 in the appendix), TPA-KV only achieves the highest average (55.03% zero-shot).
Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales.
@@
5.2 Experimental Results on FlashTPA Decoding
(a) Batch Size=1
(b) Batch Size=2
(c) Batch Size=4
(d) Batch Size=8
(e) Batch Size=16
Figure 5: Decoding time comparison of different attention mechanisms with an embedding dimension of 2048 and $d_{h}=64$. The y-axis represents $\log_{2}(\text{time})$ in seconds, and the x-axis represents $\log_{2}(\text{sequence length})$. Each subfigure corresponds to a different batch size.
This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms. We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ). It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ). While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential. Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
The evaluations were performed with batch sizes selected from $\{1,2,4,8,16\}$, model embedding dimensions ($d_{\text{model}}$) chosen from $\{1024,2048,3072\}$, and sequence lengths ranging from $2^{12}$ (4,096) to $2^{19}$ (524,288). For all experiments, the dimension per head ($d_{h}$) was fixed at 64. The ranks for TPA’s factorized components ($R_{Q},R_{K},R_{V}$) were set to $(16,1,1)$, and for GQA configurations, the number of key-value head groups was 4.
The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against
$\log_{2}(\text{sequence length})$. Lower values on the y-axis indicate faster decoding times. Results are presented in Figure 5 for an embedding dimension of 2048 (corresponding to 32 attention heads). Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix D. Figure 5 depicts these speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive and often outperforms other attention mechanisms, especially as the sequence length increases. A detailed breakdown of these comparisons across different batch sizes for the $d_{\text{model}}=2048$ case (Figure 5) is provided in Appendix D.
These findings underscore the computational efficiency of the FlashTPA decoding algorithm, particularly its favorable scaling with increasing sequence lengths.
@@ 6 Conclusion
Appendix A Expressing MHA, MQA, and GQA as Non-contextual TPA
This appendix section details how standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA).
A.1 MHA as Non-contextual TPA
Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$. More precisely, if we set $R_{Q}=R_{K}=R_{V}=h$ (the number of heads) and define the factors appropriately, MHA emerges. To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$. Focusing on $\mathbf{Q}_{t}$:
  1. (a)
    Non-contextual head factors. Define
    $\mathbf{a}^{Q}_{i}=h\mathbf{e}_{i}\in\mathbb{R}^{h},$ (A.1)
  2. (b)
    Contextual token factors. Define
    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (A.2)
    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection matrix, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ is dependent on $\mathbf{x}_{t}$.
[…]
$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (A.3)
[…]
TPA with Non-contextual A factors. More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing the token-dimension factors $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent. Then, for keys (using the original $1/R_{K}$ scaling):
and similarly for queries and values. This version, referred to as TPA (Non-contextual A), reduces per-token computations for the A factors and can be effective when head-dimension relationships are relatively stable across tokens. The KV cache for this variant would store $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ for each token, while $\mathbf{a}^{K}_{r}$ and $\mathbf{a}^{V}_{r}$ are fixed.
A.2 MQA and GQA as Non-contextual TPA
Multi-Query Attention (MQA) [45] and Grouped Query Attention (GQA) [2] (detailed in Appendix H.1 and H.2) also emerge naturally from TPA by restricting the head-dimension factors ($\mathbf{a}^{K},\mathbf{a}^{V}$) to be non-contextual and low-rank:
  • MQA as Rank-1 TPA (for K and V). In MQA, all heads share a single set of keys and values. This corresponds to TPA with $R_{K}=R_{V}=1$, where the head-dimension factors for keys and values are non-contextual vectors of all ones (or any constant non-zero vector, as it’s a shared projection). For example:
    $\mathbf{K}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
    where $\mathbf{1}_{h}\in\mathbb{R}^{h}$ is a vector of ones. This forces every head to use the same key and value vectors $\mathbf{b}^{K}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}(\mathbf{x}_{t})$ (which are effectively $\mathbf{X}\bm{W}^{K}_{\text{shared}}$ and $\mathbf{X}\bm{W}^{V}_{\text{shared}}$ from MQA’s definition). Each head $i$ would still use its distinct query $\mathbf{Q}_{t,i}$, which in TPA could be modeled with $R_{Q}=h$ and $\mathbf{a}^{Q}_{i}=\mathbf{e}_{i}$ as in the MHA case if queries are not factorized, or a different TPA structure for queries.
  • GQA as Grouped Rank-1 TPA (for K and V within groups). GQA partitions $h$ heads into $G$ groups, with each group sharing keys and values. In TPA form, this means for keys and values, the rank $R_{K}$ and $R_{V}$ would be $G$. For each group $j\in[G]$, there would be a non-contextual head-dimension factor $\mathbf{a}^{K}_{j}\in\mathbb{R}^{h}$ (and similarly $\mathbf{a}^{V}_{j}$) that acts as a "mask" or selector for the heads in that group (e.g., $\mathbf{a}^{K}_{j}$ has ones for heads in group $j$ and zeros elsewhere). Then:
    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{K}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{V}_{j}\otimes\mathbf{b}^{V}_{j}(\mathbf{x}_{t}),$
    where $\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$ is the shared key vector for group $j$. Varying $G$ from $1$ (MQA) to $h$ (MHA, if $\mathbf{b}^{K}_{j}$ are distinct per original head) shows GQA as an interpolation.
Hence, by constraining TPA’s head-dimension factors $\mathbf{a}^{K},\mathbf{a}^{V}$ to be specific constant vectors (or masks) and choosing appropriate ranks, these popular attention variants can be recovered as special cases of TPA with non-contextual A factors for keys and values. This underscores the flexibility and unifying nature of the TPA framework.
Appendix B T6 Model Architecture Details
This section provides a recap of how Tensor Product Attention (TPA) components are integrated within the Tensor ProducT ATTenTion Transformer (T6) architecture, building upon the detailed exposition in Section 3.
TPA QKV Factorization Recap. As defined in Section 3.1, for each token’s hidden-state vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, TPA projects the input into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$. The slices for the $t$-th token, $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$, are formed by sums of tensor products. The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$. For instance, for queries, with learnable weight matrices $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, these factors are:
In practice, these are often computed by merging ranks into a single linear layer followed by reshaping, as detailed in Section 3.1. The factorization for keys ($\mathbf{K}_{t}$) and values ($\mathbf{V}_{t}$) follows the same pattern using their respective ranks $R_{K},R_{V}$ and projection matrices.
Attention Step and Output Projection Recap. Once the factorized $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors are obtained for all tokens (with RoPE applied to the appropriate factors of $\mathbf{Q}$ and $\mathbf{K}$ as per Section 3.2), the attention output for each head $i\in\{1,\dots,h\}$ is computed using the scaled dot-product attention formula (Equation 3.2 from the main paper):
$\textbf{head}_{i}=\operatorname{Softmax}\Bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathbf{Q}_{i}\,(\mathbf{K}_{i})^{\top}\Bigr{)}\;\mathbf{V}_{i}.$
Finally, the outputs from all $h$ heads are concatenated and projected back to the model dimension using an output weight matrix $\bm{W}^{O}$, as shown in Equation 3.3 from the main paper:
$\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\bm{W}^{O}.$
This constitutes the output of the TPA sub-layer in a T6 block.
Appendix C Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
In TPA, the query, key, and value tensors ($\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$) are constructed as sums of $R_{Q},R_{K},R_{V}$ rank-one tensors, respectively, as detailed in Equation 3.1. These sums are typically normalized by scaling factors $s_{Q}=1/R_{Q}$, $s_{K}=1/R_{K}$, and $s_{V}=1/R_{V}$. This section explores the computational advantages of performing attention calculations directly on these factorized representations, including their scaling factors, without explicitly materializing the full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors. This approach can potentially reduce floating-point operations (FLOPs) and memory bandwidth.
C.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$